# ILCL: Inverse Logic-Constraint Learning from Temporally Constrained Demonstrations

### v1

@@ V-B Quantitative Evaluation Study with Baselines
We evaluate each methodology in both training (seen) and testing (unseen) environments. For training, we estimate a constraint and the corresponding constrained policy for each random seed. From each policy, we generate $10^{3}$ trajectories by randomizing initial conditions. In total, we produce $10^{4}$, $5\cdot 10^{3}$, and $5\cdot 10^{3}$ trajectories using $10$, $5$, and $5$ random seeds, respectively, for the three tasks presented.
For testing, we transfer the learned constraints (i.e., $10$, $5$, and $5$ constraints for each task, respectively) to novel, randomly generated environments. Creating $10$, $4$, and $4$ new environment for the three tasks, we train a constrained policy per environment. We generate $10^{3}$ trajectories varying the start or goal conditions. In total, we produce $10^{5}$, $2\cdot 10^{4}$, and $2\cdot 10^{4}$ trajectories. We then assess performance using the following metrics across all trajectories:
[…]
@@
We compare ILCL against ICL and inverse reinforcement learning (IRL) methods. The ICL methods include numeric state-action (i.e., Markovian) constraint learning approaches:
  • ICRL [11]: An inverse CRL (ICRL) method with maximum likelihood constraint inference.
  • TCL [12]: A transferable constraint learning (TCL) method by reward decomposition.
  • MTICL [6]: A multi-task ICL (MTICL) method with a two-player zero-sum game scheme (i.e., basis for ILCL).
@@
  • Option-GAIL [33]: An HIRL extension of generative adversarial imitation learning (GAIL).
  • Hier-AIRL [34]: An HIRL extension of adversarial inverse reinforcement learning (AIRL).
For evaluations, we provide all methods with identical features related to the ground-truth constraints. However, we give IRL baselines additional reward-related features to facilitate learning their reward functions.
@@ V-C Qualitative Transfer Study in Real World
[…]
@@ VI-A Quantitative Analysis through Simulations
We evaluate the constraint learning performance of ILCL against baselines across four benchmark environments with distinct ground-truth logic constraints. As shown in Fig. 5 (a), ILCL consistently outperforms baselines in learning constraints, achieving the lowest VRs in all training environments while obtaining expert-like REWs. We normalize all REW values to the expert demonstrations’ scores, ranging from $[0,1]$. Notably, in the navigation scenario with $\phi_{\text{GT1}}$, ILCL’s maximum violation rate of $32.5\%$ is $50.4\%$ lower than Hier-AIRL’s, the best performing baseline. In contrast, ICL baselines exhibit significantly higher VRs due to their inability to learn temporal structure. Although IRL baselines learn the hierarchical structure in demonstrations, the approaches frequently fail to capture either the task objective or the embedded constraint. For example, Hier-AIRL in the wiping task achieves a VR of $21\%$, lower than other baselines, but still yields low REWs due to a deficient task reward signal. In contrast, REWs exceeding $1.0$ indicate less- or un-constrained goal-seeking behaviors due to failures in constraint learning or CRL. Note that the binary nature of VRs classifies even minor, single timestep violations as full binary violations. A continuous measure is necessary to provide a more detailed assessment of constraint quality.
Figure 6: Demonstration of ILCL’s constraint transfer capability in the peg-in-shallow-hole environments. By creating real-world environments with holes tilted at $90^{\circ}$, $60^{\circ}$, and $-25^{\circ}$, we enable the Panda gripper to find a constrained policy using the learned constraints and reproduce a demonstration-like robust insertion with stable contact in each unseen environment. Demonstration videos are available in the supplementary materials including $60^{\circ}$ tilted hole insertion.
Table I: Examples of ground-truth (GT) and TLTL constraints estimated by ILCL. V, R, and T denote VR, REW, and TR, respectively. The reported statistics are the evaluation results of fully constrained trajectories generated by Logic-CRL, using the same REW and TR scales as in Fig. 5.
Env. Algo. TLTL Const. V (%) R T
@@
$<0.241^{\circ}\,\mathcal{R}\,d^{\text{peg}}_{\text{jaw}}<0.0055)\land\square d^{\text{peg}}_{\text{hole}}<0.0072)$
@@ VI-B Qualitative Demonstrations in Real World
Fig. 6 demonstrates the transferability of ILCL in novel peg-in-shallow-hole environments, where the Panda arm with a parallel jaw gripper successfully inserts a peg into unforeseen holes tilted at $90^{\circ}$, $60^{\circ}$, and $-25^{\circ}$. During the insertions, the gripper maintained stable contact between the hole and the peg. These demonstrations indicate that ILCL learns abstract logic constraints that are robustly transferable to real-world setups. For each experiment, we optimized the constrained policy for $4$ hours within $10,000$ steps of real-world interactions.
@@ VII Conclusion
We introduced an inverse logic-constraint learning (ILCL) algorithm that identifies free-form transferable TLTL constraint from demonstrations. Adopting a two-player zero-sum game, ILCL combines GA-TL-Mining—learning TLTL constraints from demonstrations without predefined templates—and Logic-CRL—optimizing policies under non-Markovian constraints. In quantitative and qualitative studies, ILCL outperforms state-of-the-art baselines, reproducing demonstration-like behaviors with fewer violations in novel environments, and generalizes to real-world peg-in-shallow-hole tasks.

### v2

@@ V-B Quantitative Evaluation Study with Baselines
We evaluate each methodology in both training (seen) and testing (unseen) environments. For training, we estimate a constraint and the corresponding constrained policy for each random seed. From each policy, we generate $10^{3}$ trajectories by randomizing initial conditions. In total, we produce $10^{4}$, $5\cdot 10^{3}$, and $5\cdot 10^{3}$ trajectories using $10$, 55, and $5$ random seeds, respectively, for the three tasks presented.
For testing, we transfer the learned constraints (i.e., $10$, 55, and $5$ constraints per task) to novel, randomly generated environments ($10$, 44, and $4$ environments per task). We train one constrained policy per environment and sample $10^{3}$ trajectories with varied starts or goals, yielding $10^{5}$, $2\cdot 10^{4}$, and $2\cdot 10^{4}$ trajectories. We then assess performance using the followings:
  • VR: the violation rate $\%$] for a ground-truth constraint $\phi_{\text{GT}}$ where any timestep violation marks the episode as violated.
@@
We compare ILCL against ICL and inverse reinforcement learning (IRL). The ICL methods include numeric state-action (i.e., Markovian) constraint learning approaches.
  • ICRL [malik2021inverse]: An inverse CRL (ICRL) method with maximum likelihood constraint inference.
  • TCL [jang2023inverse]: A transferable constraint learning (TCL) method by reward decomposition.
  • MTICL [kim2024learning]: A multi-task ICL (MTICL) method with a two-player zero-sum game scheme (i.e., basis for ILCL).
@@
  • Option-GAIL [jing2021adversarial-optiongail]: An HIRL extension of generative adversarial imitation learning (GAIL).
  • Hier-AIRL [chen2023option-hierairl]: An HIRL extension of adversarial inverse reinforcement learning (AIRL).
Figure 5: Comparison of the proposed ILCL and baseline methods in training and test environments. ‘Expert’ denotes the result of expert demonstrations without constraint violations. We normalize all REW values to the ‘Expert’ score, resulting in a range of $[0,1]$.
For evaluations, all methods receive the same features for the ground-truth constraints. IRL baselines additionally get reward-related features to learn their rewards.
@@ V-C Qualitative Transfer Study in Real World
To show the transferability, we vary the hole’s slope within $\{-25^{\circ},60^{\circ},90^{\circ}\}$. To detect slope and peg states, we use AprilTags on the gripper, hole, and peg. The learned policy applies discrete-time velocity commands, sending desired poses to a Cartesian PD controller at $1\text{\,}\mathrm{kHz}$ and RL controller at $5\text{\,}\mathrm{Hz}$.
@@ VI-A Quantitative Analysis through Simulations
We evaluate the constraint learning performance of ILCL against baselines across four benchmark environments with distinct ground-truth logic constraints. As shown in Fig. 5 (a), ILCL consistently outperforms baselines in learning constraints, achieving the lowest VRs in all training environments while obtaining expert-like REWs. We normalize all REW values to the expert demonstrations’ scores, ranging from $[0,1]$. Notably, in the navigation scenario with $\phi_{\text{GT1}}$, ILCL’s maximum violation rate of $32.5\%$ is $50.4\%$ lower than Hier-AIRL’s, the best performing baseline. In contrast, ICL baselines exhibit significantly higher VRs due to their inability to learn temporal structure. Although IRL baselines learn the hierarchical structure in demonstrations, the approaches often fail to capture either the task objective or the embedded constraint. For example, Hier-AIRL in the wiping task achieves a VR of $21\%$, lower than other baselines, but still yields low REWs due to a deficient task reward signal. In contrast, REWs exceeding $1.0$ indicate less- or un-constrained goal-seeking behaviors due to failures in constraint learning or CRL. Note that the binary nature of VRs classifies even minor, single timestep violations as full binary violations. A continuous measure is necessary to provide a more detailed assessment of constraint quality.
TABLE I: Ground-truth (GT) and the lowest VR constraints from ILCL. Reported statistics share the same scales with Fig. 5.
Env. Algo. TLTL Const. VR (%) REW TR
@@
$0.241^{\circ}\,\mathcal{R}\,d^{\text{peg}}_{\text{jaw}}<0.0055)\land\square d^{\text{peg}}_{\text{hole}}<0.0072)$
@@
Figure 6: Comparison of mean VR between ILCL and ILCL-Delay in training (left) and test (right) environments. N1, N2, W, and P indicate the evaluation tasks in the same order as in Fig. 5.
Lastly, to quantify redistribution in Logic-CRL, we compare ILCL with an ablation that removes the redistribution, which we denote as ILCL-Delay. Fig. 6 shows ILCL-Delay exhibits high VRs in sparse constraint signal tasks, such as wiping and peg-in-shallow-hole, and even reaches $100\%$ VR on the transfer experiment of navigation with $\phi_{GT2}$, highlighting the difficulty of learning new policies without redistribution. These demonstrate that constraint redistribution is crucial for handling non-Markovian temporal constraints.
@@ VI-B Qualitative Demonstrations in Real World
Fig. 7 demonstrates the transferability of ILCL in novel peg-in-shallow-hole environments, where the Panda arm with a parallel jaw gripper successfully inserts a peg into unforeseen holes tilted at $90^{\circ}$, $60^{\circ}$, and $25^{\circ}$. During the insertions, the gripper maintained stable contact between the hole and the peg. These demonstrations indicate that ILCL learns abstract logic constraints that are robustly transferable to real-world setups. For each experiment, we optimized the constrained policy for $4$ hours within $10,000$ steps of real-world interactions.
Figure 7: Demonstration of ILCL’s constraint transfer capability in the peg-in-shallow-hole environments. By creating real-world environments with holes tilted at $90^{\circ}$, $60^{\circ}$, and $25^{\circ}$, we enable the Panda gripper to find a constrained policy using the learned constraints and reproduce a demonstration-like robust insertion with stable contact in each unseen environment. Demonstration videos are available in the supplementary materials including $60^{\circ}$ tilted hole insertion.
@@ VII Conclusion
We introduced inverse logic-constraint learning (ILCL) that identifies free-form transferable TLTL constraint from demonstrations. Adopting a two-player zero-sum game, ILCL combines GA-TL-Mining—learning TLTL constraints from demonstrations without predefined templates—and Logic-CRL—optimizing policies under non-Markovian constraints. In quantitative and qualitative studies, ILCL outperforms state-of-the-art baselines, reproducing demonstration-like behaviors with fewer violations in novel environments, and generalizes to real-world peg-in-shallow-hole tasks.